OpenStudy (anonymous):

graph the function. y=4(1/5)^x

OpenStudy (solomonzelman):

\(\large\color{slate}{ y=4(1/5)^x }\) . Are you allowed to use a graphic calculator, or you need to explain the graph (sort of) ?

OpenStudy (anonymous):

graphing calculator is allowed... all i need is the graph and how u got it

OpenStudy (solomonzelman):

Use this: https://www.desmos.com/calculator/

OpenStudy (anonymous):

i have, i have the graph but i dont have a way of showing how to get it... in otherwords... we can use the calculator if it helps us to show how to graph it. ya know?

OpenStudy (solomonzelman):

So you still need to explain what the graph looks like and why?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

will start from this: our graph is: \(\large\color{slate}{ y=4(1/5)^{x} }\) when x is a big negative number: ~~~~~~~~~~~~~~~~~~~~~~~ \(\large\color{slate}{ y=4(1/5)^{-10} }\) \(\large\color{teal}{ y=4(5)^{10} }\) ~~~~~~~~~~~~~~~~~~~~~~~~ \(\large\color{slate}{ y=4(1/5)^{-100} }\) \(\large\color{teal}{ y=4(5)^{100} }\) ~~~~~~~~~~~~~~~~~~~~~~~~ \(\large\color{slate}{ y=4(1/5)^{-1000} }\) \(\large\color{teal}{ y=4(5)^{1000} }\) ~~~~~~~~~~~~~~~~~~~~~~~~

OpenStudy (solomonzelman):

see what I am doing?

OpenStudy (anonymous):

actualy... yes

OpenStudy (solomonzelman):

very good, so when x goes more and more into \(\normalsize\color{slate}{ \rm -\infty }\), the y goes where?

OpenStudy (anonymous):

opposit of the x????

OpenStudy (solomonzelman):

yes.

OpenStudy (solomonzelman):

the y becomes a very very big positive number.

OpenStudy (solomonzelman):

now, lets find the y-intercept of the function. Can you do this?

OpenStudy (solomonzelman):

given the rule that: \(\normalsize\color{royalblue}{ (b)^0=1 }\), for any real number \(\normalsize\color{royalblue}{ b }\), with an exception that \(\normalsize\color{royalblue}{ b\ne0 }\)

OpenStudy (anonymous):

i got this.... i think i am off to a good start thanks... how can medals be given???? i will give if u tell me how

OpenStudy (solomonzelman):

my connection is playing games

OpenStudy (solomonzelman):

\(\normalsize\color{royalblue}{ y=4(1/5)^x }\) plug in \(\normalsize\color{royalblue}{ 0 }\) instead of \(\normalsize\color{royalblue}{ x }\). \(\normalsize\color{royalblue}{ y=4(1/5)^0 }\). And \(\normalsize\color{royalblue}{ y=? }\)

OpenStudy (anonymous):

aha! figured it out... gave u medal thank u so much!.. the graph looks like a take off from left to right starting out beneath the x axis...

OpenStudy (solomonzelman):

but the graph will never have a y value that is zero or less.

OpenStudy (solomonzelman):

okay, as the graph \(\normalsize\color{royalblue}{ y=4(1/5)^{x} }\) has bigger x values, the smaller \(\normalsize\color{royalblue}{ 1/5 }\) will get, and that means you are multiplying 4 times a smaller number.

OpenStudy (solomonzelman):

when \(\normalsize\color{royalblue}{ x=1}\) \(\normalsize\color{royalblue}{ y=4(1/5)^{1} =4/5 }\) ~~~~~~~~~~~~~~~~~ when \(\normalsize\color{royalblue}{ x=2}\) \(\normalsize\color{royalblue}{ y=4(1/5)^{2} =4/25 }\) ~~~~~~~~~~~~~~~~~ when \(\normalsize\color{royalblue}{ x=3}\) \(\normalsize\color{royalblue}{ y=4(1/5)^{3} =4/125 }\) ~~~~~~~~~~~~~~~~~ when \(\normalsize\color{royalblue}{ x=4}\) \(\normalsize\color{royalblue}{ y=4(1/5)^{4} =4/625 }\) ~~~~~~~~~~~~~~~~~ when \(\normalsize\color{royalblue}{ x=5}\) \(\normalsize\color{royalblue}{ y=4(1/5)^{5} =4/3125 }\)

OpenStudy (solomonzelman):

so it will be hitting zero closer and closer, so it should look like